OpenStudy (anonymous):

Find tan of angle B. Write your answer as a simplified fraction. Use the triangle below to answer this question.

OpenStudy (anonymous):

triangle?

OpenStudy (anonymous):



OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

How Did you do that?

OpenStudy (anonymous):

I got 3/4

OpenStudy (anonymous):

which answer is correct? mine one or yours?

OpenStudy (mayankdevnani):

@CynthiaJune94 do you know tan theta formula?

OpenStudy (anonymous):

Tan = Opposite/Adjacent

OpenStudy (anonymous):

tan = perp/base and in your case tan = 15/20

OpenStudy (anonymous):

Yea and when i divided 15/20 i got 3/4

OpenStudy (mayankdevnani):

good!

OpenStudy (anonymous):

but you said it was 1/4

OpenStudy (anonymous):

yeah sorry it will be 3/4..:P

OpenStudy (mayankdevnani):

so what's your problem!! answer is correct!! @CynthiaJune94

OpenStudy (anonymous):

Okay thanks now what about the Sec of angle B Sec= Hypotenuse/Adjacent Angle B=25/20 or 1 1/4 Is this correct too? And i dont really get 'Trig. It is really hard for me.

OpenStudy (anonymous):

trig is very interesting... believe me.. :)

OpenStudy (anonymous):

Are you in Trig @sehanhasan?

OpenStudy (anonymous):

may be i do not remember..

OpenStudy (anonymous):

send me the link.. i will be there

OpenStudy (anonymous):

link to what?

OpenStudy (anonymous):

what do u mean by Are you in Trig? u mean in Trig group..? :S

OpenStudy (mayankdevnani):

no its 5/4

OpenStudy (mayankdevnani):

ok @CynthiaJune94

OpenStudy (anonymous):

in a class for trig... and that's the same thing and it said in simplest fraction

OpenStudy (mayankdevnani):

yaaa, 25/20=5/4=sec of angle B

OpenStudy (anonymous):

alrite

OpenStudy (anonymous):

why can't it be 1 1/4?

OpenStudy (mayankdevnani):

hey!!! 25/20 write in simplest form

OpenStudy (anonymous):

Okay thanks sorry i just didn't understand Now if I switch to angle A and have to find the Cot Cot=Adjacent/Opposite Angle A=15/20= 3/4 Would this be correct?

OpenStudy (mayankdevnani):

yaaa, you are 100% correct

OpenStudy (anonymous):

So Csc of Angle A Csc=Hypotenuse/Opposite Angle A=25/20=5/4 Correct?

OpenStudy (mayankdevnani):

yaaa correct

OpenStudy (anonymous):

And Cos of Angle A Cos=Adjacent/Hypotenuse Angle A=20/25=4/5 Also Correct?

OpenStudy (mayankdevnani):

wrong!!!

OpenStudy (mayankdevnani):

Cos=Adjacent/Hypotenuse Angle A=15/25=3/5

OpenStudy (anonymous):

Oh okay thank! Simple mistake sorry. Okay thanks I think i got the hang of this now!!! Thanks for all your help!!!! You helped a lot!